# Evolution of interior entropy of a loop quantum-corrected black hole pierced by a cosmic string

### v1

@@ I Introduction
On the other hand, the authors of Lewandowski:2022zce apply loop quantum gravity(LQG) method and obtain a loop quantum-corrected Schwarzschild black hole who contains quantum gravity effects in the metric directly via quantum Oppenheimer-Snyder model. The effects of quantum gravity of the obtained black hole could hidden in the observable hole’s shadowYang:2022btw . Besides, the hole’s singularity will be remove and there exists a minimal mass of the hole at Planck scale which means that there will be a stable remnant at the end of Hawking radiation for the black hole. Furthermore, Han:2023wxg finds a geometry of black to white hole transition for the loop quantum-corrected Schwarzschild black hole and Han:2024rqb further calculates the spin foam amplitude of this quantum tunneling transition in LQG. These works implying that the black hole information may be eventually either store in a remnant or release through a white hole if nature admits it. If so, the concept of interior entropy could provide a choice to explain where the information goes before the black hole shrinks to a remnant or appearing white holeRovelli:2024sjl . Theoretically, if interior entropy could serve as a new type of entropy which is related to informationPage:1993wv ; Rovelli:2017mzl , the total variation between this type of entropy and Bekenstein-Hawking entropy is expected to satisfy the second law of thermodynamics during Hawking radiation.
In addition, quantum field theory predicts that it may exist topology defects such as cosmic stringsHindmarsh:1994re ; Copeland:2009ga due to occurring of phase transition which induced by spontaneous symmetry breaking in early universe. Cosmic strings would survive in the universe in the form of network which consists of infinite cosmic strings and cosmic string loops. Also, cosmic strings can give rise to some observable effects. For example, when the string network evolves in the spacetime, cosmic strings will radiate gravitational wavesVachaspati:1984gt ; Allen:1991bk ; Damour:2004kw , perhaps form black holesGarriga:1992nm , and could result in gravitational lensing phenomenonVilenkin:1984ea . A recent work in Aurrekoetxea:2023vtp takes numerical relativity methodHelfer:2018qgv ; Aurrekoetxea:2020tuw to investigate GW190521 event of gravitational wave which obtained from the observation results in the third LIGO-Virgo-KAGRA run, and finds that this signal likely originates from a collapsing cosmic string loop. Another phenomenon may arise in astrophysics is that cosmic strings connect to black holes and form topology defect black holesAryal:1986sz . It is still unknown that how do cosmic strings influence black hole’s interior entropy as well as its relation to Bekenstein-Hawking entropy during black hole evaporation once this kind of topology defect is considered as all the previously works do not involve this aspect.
@@ III Interior entropy and its relation to Bekenstein-Hawking entropy
$T_{H}=\frac{\kappa}{2\pi}=\frac{1}{4\pi}\frac{df(r_{+})}{dr}=\frac{M}{2\pi r_{+}^{2}}(1-\frac{2\alpha M}{r_{+}^{3}}),$ (16)
@@
and it is obvious that the evolution relation between this two types of entropy is depend on the string tension $\mu$ and the quantum gravity effects characterized by $\alpha$. Also, notice here that the corresponding evolution relation of a Schwarzschild black hole pierced by a cosmic string can be obtained in the absence of quantum effects when letting $\beta\to 1$ (i.e., $\alpha\to 0$ and $r_{+}\to 2M$) such that

### v2

@@ I Introduction
On the other hand, the authors of Lewandowski:2022zce apply loop quantum gravity (LQG) method and obtain a loop quantum-corrected Schwarzschild black hole who contains quantum gravity effects in the metric directly via quantum Oppenheimer-Snyder model. The effects of quantum gravity of the obtained black hole could hide in the observable hole’s shadowYang:2022btw . Besides, the hole’s singularity will be removed and there exists a minimal mass of the hole at Planck scale which means that there will be a stable remnant at the end of Hawking radiation for the black hole. Furthermore, Han:2023wxg finds a geometry of black to white hole transition for the loop quantum-corrected Schwarzschild black hole and Han:2024rqb further calculates the spin foam amplitude of this quantum tunneling transition in LQG. These works implying that the black hole information may be eventually either store in a remnant or release through a white hole if nature admits it. If so, the concept of interior entropy could provide a choice to explain where the information goes before the black hole shrinks to a remnant or appearing white holeRovelli:2024sjl . Theoretically, if interior entropy could serve as a new type of entropy which is related to informationPage:1993wv ; Rovelli:2017mzl , the total variation between this type of entropy and Bekenstein-Hawking entropy is expected to satisfy the second law of thermodynamics during Hawking radiation.
In addition, quantum field theory predicts that it may exist topology defects such as cosmic stringsHindmarsh:1994re ; Copeland:2009ga due to occurring of phase transition which induced by spontaneous symmetry breaking in early universe. Cosmic strings would survive in the universe in the form of network which consists of infinite cosmic strings and cosmic string loops. Also, cosmic strings can give rise to some observable effects. For example, when the string network evolves in the spacetime, cosmic strings will radiate gravitational wavesVachaspati:1984gt ; Allen:1991bk ; Damour:2004kw , perhaps form black holesGarriga:1992nm , and could result in gravitational lensing phenomenonVilenkin:1984ea . A recent work in Aurrekoetxea:2023vtp takes numerical relativity methodHelfer:2018qgv ; Aurrekoetxea:2020tuw to investigate GW190521 event of gravitational wave which obtained from the observation results in the third LIGO-Virgo-KAGRA run, and finds that cosmic string loops that collapse to black holes could potentially mimic high-mass binary black hole mergers if only the ringdown is observed. Another phenomenon may arise in astrophysics is that cosmic strings connect to black holes and form topology defect black holesAryal:1986sz . It is still unknown that how do cosmic strings influence black hole’s interior entropy as well as its relation to Bekenstein-Hawking entropy during black hole evaporation once this kind of topology defect is considered as all the previously works do not involve this aspect.
@@ III Interior entropy and its relation to Bekenstein-Hawking entropy
$T_{H}=\frac{\kappa}{2\pi}=\frac{1}{4\pi}\frac{df(r_{+})}{dr}=\frac{M}{2\pi r_{+}^{2}}(1-\frac{2\alpha M}{r_{+}^{3}}).$ (16)
@@
and it is obvious that the evolution relation between this two types of entropy depends on the string tension $\mu$ and the quantum gravity effects characterized by $\alpha$. Also, notice here that the corresponding evolution relation of a Schwarzschild black hole pierced by a cosmic string can be obtained in the absence of quantum effects when letting $\beta\to 1$ (i.e., $\alpha\to 0$ and $r_{+}\to 2M$) such that